# ChemGen: Code Generation for Multispecies Chemical Kinetics in Physics Simulations

### v1

@@ Abstract
Developer’s repository link: https://github.com/drryjoh/chemgen/
@@ 1 Introduction
Modeling of chemical reactions is critical to many areas of computational physics. One of the most common examples of this coupling is in Computational Fluid Dynamics (CFD) [1, 2, 3], which is used to simulate fluid behavior in a wide variety of applications. The governing equations, typically the Navier-Stokes equations, are discretized in space using approaches such as finite volume [4, 5, 6], finite difference [7], and finite element schemes [8, 9]. CFD relies on large volumes of data to describe the fluid; in particular, the fluid state variables stored at all grid points in the computational domain are known as the degrees of freedom (DoF).
Developing robust CFD software is difficult due to computational challenges such as efficient DoF communication in parallel, adaptive physical modeling, and managing complex geometries. Additionally, there are significant numerical issues, such as preserving stability [10, 11, 12, 13, 14, 15, 16, 17] which is coupled with the choice of formulation of the conservation laws and time integration strategy. There is no single CFD software that simultaneously addresses all the above difficulties for all fluid dynamics applications, necessitating continued research into accelerating simulations while achieving stable and accurate fluid modeling. As a result, a broad family of CFD codes has emerged, ranging from experimental in-house tools used for demonstrating new technologies to established commercial software suites employed in production-level engineering.
A typical three-dimensional simulation of compressible, non-reacting flow involves five conserved variables: energy, mass, and three momentum components, such that $n_{y}=5$, where $n_{y}$ is the total state size. In many cases, existing non-reacting CFD software built around these five variables can be extended to chemically reacting flows by expanding the state to include multiple chemical species and incorporating $n_{r}$ chemical reactions. The state size then becomes $n_{y}=n_{s}+4$ , where $n_{s}$ represents the number of chemical species [^0] The original five state variables are reduced by one since total density can be determined from the species concentrations (or partial densities).. Detailed chemical models may involve hundreds of species and thousands of reactions, often with $n_{r}\approx 5n_{s}$ [18]. The number of species increases with the complexity of the fuel; for instance, the hydrogen model in [19, 17] consists of 10 chemical species and 17 reactions, while the FFCM-2 model for simple hydrocarbons (up to four carbons per chemical species) comprises 96 species and 1054 reactions [20].
@@
    Adapt existing legacy software, such as Chemkin [21], for the target CFD code.
@@
    Use the more widely adopted, open-source, and actively maintained Cantera [22], which can serve as a third-party library for codes written in C, C++ , and Fortran, as well as in interpreted languages such as Python and MATLAB.
@@
Approaches 1 and 2 are reasonable choices but present certain challenges. With approach 1, the existing legacy routines may not easily adapt to the specific requirements of the targeted CFD software, limiting scalability and complicating code maintenance. Regarding approach 2, although the package is excellent for property calculations , it can only act as a third-party library, which may lead to limited control over software functionality. Additionally, relying on a third-party library(e.g. Cantera and Mutation++ [23]) may hinder compatibility with targeted architectures that developers cannot directly access. Using approach 3 can also introduce challenges; it often requires extensive work by developers, as the necessary models would need to be built from scratch (further described in Section 2). This work belongs to approach 4 and delivers a tool to generate source code that is compatible with the existing ecosystem of a target software package.
Code generation is a well-established concept, with a history in chemistry applications. There are two publicly available code generators focused on gas-phase combustion: PyJac [24] and Pyrometheus [25]. PyJac specializes in generating analytical source term Jacobians, which can be used to accelerate the computation of chemical species evolution. Pyrometheus generates either C++ header files or Python files for gas-phase combustion calculations, incorporating several on-the-fly optimizations to enhance scalability for high-performance computing in CFD. Additionally, KinetiX [26] is a software toolkit that generates chemical source terms as well as thermodynamic and mixture-averaged transport properties tailored to specific hardware targets, such as CPU or GPU architectures.
@@
In this work, we introduce ChemGen, a software package that leverages code generation to integrate multispecies thermodynamics and chemical kinetics into C++ based codes. ChemGen introduces decorators, discussed in detail in Section (4), to enable flexible C++ code generation tailored to existing software ecosystems. ChemGen currently focuses on thermodynamic properties and chemical source terms, as well as their analytical derivatives. The analytical derivatives facilitate Jacobian assembly for implicit time integration strategies appropriate for CFD. In addition, a number of time integrators, linear solvers, and preconditioners is provided for convenience.
@@
This paper is organized as follows. We begin with an overview of the chemically reacting flow conservation equations used in CFD, highlighting where chemistry fits within these formulations, including details on multi-species thermodynamics and chemical reaction rates. Next, we discuss time integration strategies, focusing on operator splitting approaches that reduce the problem to a system of ordinary differential equations for chemical state evolution, as well as fully coupled formulations. We then introduce the code generation process, emphasizing how malleability is achieved through the use of decorators. Finally, we present results verifying the source term evaluation, Jacobian computation, and time integration. We demonstrate ChemGen’s performance by replacing OpenFOAM’s existing ODE-based chemical integrator in a detonation simulation and achieve faster time-to-solution due to ChemGen’s specialized, generated source code.
@@ 2 Governing equations and thermodynamics
where $d$ is the spatial dimension, $n_{s}$ is the number of species (which yields a state size $n_{y}=d+n_{s}+1$), $\rho$ is density, $v=\left(v_{1},\ldots,v_{d}\right)$ is the velocity vector, $e_{t}$ is the mass-specific total energy, and $C=\left(C_{1},\ldots,C_{n_{s}}\right)$ is the vector of molar species concentrations. The density is computed from the species concentrations as
@@
where $\rho_{i}$ is the partial density and $W_{i}$ is the molecular mass of the $i$th species. The mass fraction of the $i$th species is defined as
@@
Beyond the chemical state values, several additional parameters are required in kinetics. The first is the pressure, obtained from the ideal gas law,
@@
with $R=8314.4626\,\mathrm{J\cdot kmol^{-1}K^{-1}}$ the universal gas constant. Another common parameter, used to reduce repeated logarithm evaluations in reaction optimizations, is $\ln T$. Finally, the mixture concentration,
@@
is required for evaluating many reaction rates, including third-body reactions with default efficiencies. In what follows, we assume that the elements of $y_{c}$, together with $\ln T$, $p$, and $M_{c}$, are available throughout ChemGen. For brevity, functional dependencies are omitted.
@@ 2.1 Thermodynamics
In this section, we define the relevant thermodynamic quantities for the $i$th species assuming a thermally perfect gas. The mass-specific internal energy is given by [27]
@@
where $h_{i}$ is the mass-specific enthalpy, $R_{i}=R/W_{i}$, and $T_{\mathrm{ref}}=298\text{ K}$ is the reference temperature. The mass-specific enthalpy, $h_{i}$, is given as
@@
where $h_{\mathrm{ref},i}$ is the reference-state enthalpy of formation, and $c_{p,i}$ is the mass-specific heat capacity at constant pressure. For differentiability and computational efficiency, ChemGen fits the polynomial for $c_{p,i}$ to arbitrary order. As such, the value of $c_{p,i}$ is computed from an $n_{p}$-order polynomial as
@@
which is based on, but not equivalent, to the NASA n-coefficient polynomial parametrization [28, 29]. Substituting Equation (9) into Equation (8), we evaluate the mass-specific enthalpy as
@@
The mass-specific thermodynamic entropy of the mixture is defined as
@@
where $s_{\mathrm{ref},i}^{o}$ is the species entropy of formation from the NASA n-coefficient polynomial parametrization evaluated at $T=298$ K, and $p_{ref}$ is the reference pressure (usually $p_{\mathrm{ref}}=1\text{ atm}$), $s_{i}^{o}$ denotes the mass-specific entropy at the reference pressure, and $p_{i}=C_{i}RT$ is the partial pressure. The term, $R\log\frac{p_{i}}{p_{\mathrm{ref}}}$, accounts for the effect of pressure on entropy and does not depend on thermodynamic polynomial fits. Using Equation (9), the mass-specific standard entropy, $s_{i}^{o}$, is computed as
@@
Finally, the Gibbs free energy of the $i$th species is
@@
Note, the Gibbs free energy, like entropy, has a pressure dependency that is calculated independent of these thermodynamic fits. Table (1) summarizes the thermodynamic relationships used by ChemGen.
Thermodynamic Quantity of the $i$th Species Fit Coefficient Description
Mass-specific heat at constant pressure, $c_{p,i}$ $c_{p,i}=\sum_{k=0}^{n_{p}}a_{ik}T^{k}$ Re-fit from existing data [28, 29]
@@
The Jacobian of the $\rho u\left(C,T\right)$with respect to temperature is given by
@@
where $c_{v}$ is the specific heat at constant volume, which is related to $c_{p}$ via
@@
where $n$ is the number of iterations and $T_{0}$ is the initial guess for temperature, e.g., $T_{0}=500\textrm{K}$. In ChemGen, we set $n=5$, which, in our experience, is sufficient to be within $1\times 10^{-8}$ relative error for a variety of models at different conditions.
@@ 2.2 Chemical reaction rates and their ChemGen representations
The source term in Equation (1) is a function of the fluid state, written as [30]
@@
where $\omega_{i}$ is the production rate of the $i$th specie, which satisfies mass conservation:
@@
The production rate is computed as
$\omega_{i}=\sum_{j=1}^{n_{r}}\nu_{ij}q_{j},$
where $n_{r}$ is the number of reactions, $\nu_{ij}=\nu_{ij}^{r}-\nu_{ij}^{f}$ is the difference between the reverse ($\nu_{ij}^{r}$) and the forward stoichiometric coefficients ($\nu_{ij}^{f}$), and $q_{j}$ is the rate of progress of the $j$th reaction, computed as
@@
where $k_{j}^{f}$ and $k_{j}^{r}$ are the forward and reverse rate constants, respectively, of the $j$th reaction. The forward and reverse rate constants are related via the equilibrium constant
@@
with $\Delta G_{j}^{\prime}$ denoting the change in reference-state Gibbs free energy for the $j$th reaction, given as
@@ 2.2.1 Arrhenius reactions
where $A_{j}$ is the prefactor, $b_{j}$ is the temperature exponent, and $E_{j}$ is the activation energy [27, 30]. In ChemGen, an optimization is made to again avoid the power function. Generally, $b_{j}$ is not an integer, however, since the exponential function is unavoidable, we can utilize the precomputed $\ln T$ and insert an additional term in the exponent
@@
2.2.2 Three-body reactions
Third-body reactions require the inclusion of the mixture concentration, which represents the collective effect of all species acting as a third body. Dissociation and recombination reactions are often of this type. The rate of progress is scaled by a prefactor as [30]
@@ 2.2.2 Three-body reactions
which is still attributed to the second term in Rquation (37) via the relationship of $k_{j}^{f}$ to $k_{j}^{r}$ from Equation (35). An additional optimization is made during the code generation process for the $\alpha_{ij}$ terms. The established convention is that the efficiencies always default $\alpha_{ij}=1$. Therefore, for any species without an explicitly prescribed efficiency, the $\left(\alpha_{ij}-1\right)C_{i}$ term is removed during code generation. This avoids possibly repetitive work to determine the efficiency from a sum of products of all species to a sum of products of all species with $\alpha_{ij}\neq 1$ by utilizing the precomputed $M_{c}$.
2.2.3 Fall-off reactions
Unimolecular/recombination fall-off reactions incorporate a dependence on pressure. In general, this model predicts an increase in the reaction rate with increasing pressure. Given Arrhenius-type low-pressure and high-pressure limits for the rate functions ($k_{A,0,j}$ and $k_{A,\infty,j}$, respectively with their own $A_{j}$, $b_{j}$, and $E_{j}$ constants), $k_{j}^{f}$ is computed as

### v3

# ChemGen: Code Generation for Multispecies Chemically Reacting Flow Simulations

@@ Abstract
Developer’s repository link: https://github.com/drryjoh/chemgen
@@ 1 Introduction
Modeling of chemical reactions is critical to many areas of computational physics. One of the most common examples is in modeling the chemistry-turbulence coupling in Computational Fluid Dynamics (CFD) [And95, Fer02, Hir07], which is used to simulate fluid behavior in a wide variety of reacting flows and combustion applications. The governing equations, typically the Navier-Stokes equations, are discretized in space using approaches such as finite volume [Pat80, Ver07, Mou16], finite difference [LeV07], and finite element schemes [Don03, Zie05]. CFD relies on large volumes of data to describe the fluid; in particular, the fluid state variables stored at all grid points in the computational domain are known as the degrees of freedom (DoF).
Developing robust CFD software is often challenging due to computational challenges such as efficient DoF communication in parallel, adaptive physical modeling, and managing complex geometries. Additionally, there are significant numerical issues, such as preserving stability [Kar94, Abg96, Bil03, Lv15, Lv17, Joh20, Chi22, Chi22_2] which is coupled with the choice of formulation of the conservation laws and time integration strategy. There is no single CFD software that simultaneously addresses all the above difficulties for all fluid dynamics applications, necessitating continued research into accelerating simulations while achieving stable and accurate fluid modeling. As a result, a broad family of CFD codes has emerged, ranging from experimental in-house tools used for demonstrating new technologies to established commercial software suites employed in production-level engineering.
A typical three-dimensional simulation of compressible, non-reacting flow involves five conserved variables: energy, mass, and three momentum components, such that $n_{y}=5$, where $n_{y}$ is the total state size. In many cases, existing non-reacting CFD software built around these five variables can be extended to chemically reacting flows by expanding the state to include multiple chemical species and incorporating $n_{r}$ chemical reactions. The state size then becomes $n_{y}=n_{s}+4$ , where $n_{s}$ represents the number of chemical species [^0] The original five state variables are reduced by one since total density can be determined from the species concentrations (or partial densities).. Detailed chemical models may involve hundreds of species and thousands of reactions, often with $n_{r}\sim 5n_{s}$ [Lu09]. The number of species increases with the complexity of the fuel; for instance, the hydrogen model in [Wes82, Chi22_2] consists of 10 chemical species and 17 reactions, while the FFCM-2 model for simple hydrocarbons (up to four carbons per chemical species) comprises 96 species and 1054 reactions [Zha23].
@@
    Adapt existing legacy software, such as Chemkin [chemkin89], for the target CFD code.
@@
    Use the more widely adopted, open-source, and actively maintained Cantera [cantera], which can serve as a third-party library for codes written in C, C++ , and Fortran, as well as in interpreted languages such as Python and MATLAB.
@@
Approaches 1 and 2 are reasonable choices but present certain challenges. With approach 1, the existing legacy routines may not easily adapt to the specific requirements of the targeted CFD software, limiting scalability and complicating code maintenance. Regarding approach 2, although the package is excellent for property calculations , it can only act as a third-party library, which may lead to limited control over software functionality. Additionally, relying on a third-party library(e.g. Cantera and Mutation++ [Sco20]) may hinder compatibility with targeted architectures that developers cannot directly access. Using approach 3 can also introduce challenges; it often requires extensive work by developers, as the necessary models would need to be built from scratch (further described in Section 2). This work belongs to approach 4 and delivers a tool to generate source code that is compatible with the existing ecosystem of a target software package.
Code generation has a long history in chemical kinetics. Three public tools focus on gas-phase combustion: PyJac [Nie17], Pyrometheus [Cis25arxiv], and KinetiX [Dan25]. PyJac generates analytical source-term Jacobians to accelerate species evolution. Pyrometheus produces C++ headers or Python modules for gas-phase kinetics with on-the-fly optimizations for scalable CFD. KinetiX generates source terms, thermodynamic properties, and mixture-averaged transport models for CPU and GPU targets. These tools broaden access to high-performance chemistry through code generation, but, unfortunately, none of them provide a single framework that unifies source-term generation, Jacobian construction within a differentiable code base, and stand-alone time integration strategies, all of which are central to chemically reacting flow simulations.
@@
In this work, we introduce ChemGen, a software package that leverages code generation to integrate multispecies thermodynamics, chemical kinetics, and time integration strategies into C++ based codes. ChemGen introduces decorators, discussed in detail in Section (4), to enable flexible C++ code generation tailored to existing software ecosystems. ChemGen currently focuses on thermodynamic properties and chemical source terms, as well as their analytical derivatives. The analytical derivatives facilitate Jacobian assembly for implicit time integration strategies appropriate for CFD. In addition, a number of time integrators, linear solvers, and preconditioners is provided for convenience.
@@
This paper is organized as follows. We begin with an overview of the chemically reacting flow conservation equations used in CFD, highlighting where chemistry fits within these formulations, including details on multi-species thermodynamics and chemical reaction rates. Next, we discuss time integration strategies, focusing on operator splitting approaches that reduce the problem to a system of ordinary differential equations for chemical state evolution, as well as fully coupled formulations. We then introduce the code generation process, emphasizing how malleability is achieved through the use of decorators. Finally, we present results verifying the source term evaluation, Jacobian computation, and time integration. We demonstrate ChemGen’s performance by replacing OpenFOAM’s existing ODE-based chemical integrator in a detonation simulation and achieve significantly faster time-to-solution due to ChemGen’s specialized, generated source code. Overall, we show that ChemGen provides a path forward for improved malleability and computational efficiency compared to existing chemical kinetic solvers used in high-level CFD applications.
@@ 2 Governing equations and thermodynamics
where $d$ is the spatial dimension, $n_{s}$ is the number of species (which yields a state size $n_{y}=d+n_{s}+1$), $\rho$ is density, $v=\left(v_{1},\ldots,v_{d}\right)$ is the velocity vector, $e_{t}$ is the mass-specific total energy, and $C=\left(C_{1},\ldots,C_{n_{s}}\right)$ is the vector of species molar concentrations. The density is computed from the species concentrations as
@@
where $\rho_{i}$ is the partial density and $W_{i}$ is the molecular mass of the $i^{th}$ species. The mass fraction of the $i^{th}$ species is defined as
@@
Beyond the chemical state values, several additional parameters are required. The first is the pressure, which may be obtained from the ideal gas law,
@@
with $R$ is the universal gas constant. Nonideal gas treatment is possible for the equation of state and relevant thermodynamic properties, but this has not been attempted in the current version of ChemGen. Another common parameter, used to reduce repeated logarithm evaluations in reaction optimizations, is $\ln T$. Finally, the mixture concentration,
@@
is required for evaluating pressure-dependent reaction rates. In what follows, we assume that the elements of $y_{c}$, together with $\ln T$, $p$, and $M_{c}$, are available throughout ChemGen. For brevity, functional dependencies are omitted.
@@ 2.1 Thermodynamics
In this section, we define the relevant thermodynamic quantities for the $i^{th}$ species. The mass-specific internal energy is given by [Gio99]
@@
where $h_{i}$ is the mass-specific enthalpy and $R_{i}=R/W_{i}$. The mass-specific enthalpy, $h_{i}$, is given as
@@
where $T_{\mathrm{ref}}=298\text{ K}$ is the reference temperature, $h_{\mathrm{ref},i}$ is the standard enthalpy of formation at $T_{\mathrm{ref}}$, and $c_{p,i}$ is the mass-specific heat capacity at constant pressure. For differentiability and computational efficiency, ChemGen fits the polynomial for $c_{p,i}$ to arbitrary order. As such, the value of $c_{p,i}$ is computed from an $n_{p}$-order polynomial as
@@
which is based on, but not equivalent, to the NASA n-coefficient polynomial parametrization [Mcb93, Mcb02]. Substituting Equation (9) into Equation (8), we evaluate the mass-specific enthalpy as
@@
The mass-specific entropy of the mixture is defined as
@@
where $s_{\mathrm{ref},i}^{o}$ is the species entropy of formation from the NASA n-coefficient polynomial parametrization evaluated at $T=298$ K, and $p_{ref}$ is the reference pressure (usually $p_{\mathrm{ref}}=1\text{ atm}$), $s_{i}^{o}$ denotes the mass-specific entropy at the reference pressure, and $p_{i}=C_{i}RT$ is the partial pressure. The term, $R\log\frac{p_{i}}{p_{\mathrm{ref}}}$, derives from the Gibbs equation, which accounts for the effect of pressure on entropy, and does not depend on thermodynamic polynomial fits. Using Equation (9), the mass-specific standard entropy, $s_{i}^{o}$, is computed as
@@
Finally, the Gibbs free energy of the $i^{th}$ species is
@@
Note, the Gibbs free energy, like entropy, has a pressure dependency that is calculated independent of the thermodynamic fits. Table (1) summarizes the thermodynamic relationships used by ChemGen.
Thermodynamic Quantity of the $i^{th}$ Species Fit Coefficient Description
Mass-specific heat at constant pressure, $c_{p,i}$ $c_{p,i}=\sum_{k=0}^{n_{p}}a_{ik}T^{k}$ Re-fit from existing data [Mcb93, Mcb02]
@@
The Jacobian of $\rho u\left(C,T\right)$ with respect to temperature is given by
@@
where $c_{v}$ is the constant-volume specific heat,
@@
where $n$ is the number of iterations and $T_{0}$ is the initial guess for temperature. In ChemGen, we set $n=5$, which, in our experience, is sufficient to be within $1\times 10^{-8}$ relative error for a variety of models at different conditions.
@@ 2.2 Chemical reaction rates and their ChemGen representations
The source term in Equation (1) is a function of the fluid state, written as [Kee96]
@@
where $\omega_{i}$ is the production rate of the $i^{th}$ specie, which satisfies mass conservation:
@@
For an arbitrary reaction model written as
$\sum_{i=1}^{n_{s}}\nu_{ij}^{f}A_{i}\rightleftharpoons\sum_{i=1}^{n_{s}}\nu_{ij}^{r}A_{i}\ (j=1,...n_{r}),$ (34)
where $A_{i}$ refers to the $i^{th}$ species $\nu_{ij}$ refers the stoichiometric coefficient of species $i$ in reaction $j$ with superscript $f$ and , the production rate is computed as
$\omega_{i}=\sum_{j=1}^{n_{r}}\nu_{ij}q_{j}.$
In the above equation, $n_{r}$ is the number of reactions, $\nu_{ij}=\nu_{ij}^{r}-\nu_{ij}^{f}$ is the difference between the reverse ($\nu_{ij}^{r}$) and the forward stoichiometric coefficients ($\nu_{ij}^{f}$), and $q_{j}$ is the rate of progress of the $j^{th}$ reaction, computed as
@@
where $k_{j}^{f}$ and $k_{j}^{r}$ are the forward and reverse rate constants, respectively, of the $j^{th}$ reaction. The forward and reverse rate constants are related via the equilibrium constant
@@
with $\Delta G_{j}^{\prime}$ denoting the change in reference-state Gibbs free energy for the $j^{th}$ reaction, given as
@@ 2.2.1 Arrhenius rate expression
where $A_{j}$ is the prefactor, $b_{j}$ is the temperature exponent, and $E_{j}$ is an quasi-activation energy [Gio99, Kee96]. In ChemGen, an optimization is made to again avoid the power function. Generally, $b_{j}$ is not an integer, however, since the exponential function is unavoidable, we can utilize the precomputed $\ln T$ and insert an additional term in the exponent
@@
2.2.2 Pressure-dependent reactions
Evaluation of rates of pressure-dependent reactions requires the use of the mixture concentration, which represents the collective effect of all species acting as a third body in collisional activation or deactivation. Dissociation and recombination reactions are often of this type. The rate of progress is scaled by a prefactor as [Kee96]
@@ 2.2.2 Pressure-dependent reactions
which is still attributed to the second term in Rquation (38) via the relationship of $k_{j}^{f}$ to $k_{j}^{r}$ from Equation (36). An additional optimization is made during the code generation process for the $\alpha_{ij}$ terms. The established convention is that the efficiencies always default $\alpha_{ij}=1$. Therefore, for any species without an explicitly prescribed third-body or Chaperone efficiency efficiency, the $\left(\alpha_{ij}-1\right)C_{i}$ term is removed during code generation. This avoids possibly repetitive work to determine the Chaperon efficiency from a sum of products of all species to a sum of products of all species with $\alpha_{ij}\neq 1$ by utilizing the precomputed $M_{c}$.
2.2.3 Pressure fall-off effect
Unimolecular/recombination reactions in the pressure fall-off region require special consideration. Given Arrhenius-type low-pressure and high-pressure limits for the rate functions ($k_{A,0,j}$ and $k_{A,\infty,j}$, respectively with their own $A_{j}$, $b_{j}$, and $E_{j}$ constants), $k_{j}^{f}$ is computed as